A point of charge 10×10^-6C has a potential of 4×10^4V from a point 2mm from it. Calculate the potential due to the charge when the distance is now 16mm.

1
Expert's answer
2022-02-11T08:48:38-0500

Given:

"V_1=4*10^4\:\rm V"

"r_1=2*10^{-3}\:\rm m"

"r_2=16*10^{-3}\:\rm m"


The potential of electric field of point charge

"V=k\frac{q}{r}\sim\frac{1}{r}"

Hence

"V_2=V_1\frac{r_1}{r_2}=4*10^4*\frac{0.002}{0.016}=5000\:\rm V"